# On the Generalization of SFT: A Reinforcement Learning Perspective with Reward Rectification

### v1

@@ 4.1.2 Main Results
DFT demonstrates generalization and robustness, especially on challenging benchmarks where standard SFT yields minimal or even negative impact. For instance, on Olympiad Bench, SFT degrades performance for Qwen2.5-Math-1.5B, dropping accuracy from 15.88 to 12.63, while DFT  boosts it to 27.08, +11.20 point improvement over base model. On AIME24, SFT reduces accuracy for Qwen2.5-Math-7B by 4.20 points (from 6.68 to 2.48), whereas DFT improves performance to 8.56, achieving a +1.88 point gain over the base model despite the difficulty of the benchmark. A similar trend is observed on AMC23. SFT reduces the performance of Qwen2.5-Math-1.5B from 19.38 to 18.75, while DFT raises it to 38.13, a +18.75 point gain over base. For Qwen2.5-Math-7B, SFT yields only a marginal improvement (+1.86), whereas DFT achieves a +17.04 point gain. These results underscore that DFT not only scales more effectively across models of varying capacities, but also exhibits greater resilience on difficult reasoning tasks where traditional SFT struggles. This highlights its potential as a more robust fine-tuning paradigm to enhance mathematical reasoning capabilities in LLMs.
DFT exhibits better learning efficiency and faster convergence characteristics.
Figure 1 reveals clear differences in learning dynamics between DFT and standard SFT on Qwen2.5-Math-1.5B across all math reasoning benchmarks. Compared to SFT, our method demonstrates three distinct advantages: (1) Faster convergence, achieving peak performance within the first 120 training steps on most benchmarks; (2) Better early-stage performance, with DFT already outperforming best final accuracy of SFT within the first 10–20 steps; and (3) Higher sample efficiency, consistently requiring fewer updates to reach relatively optimal results. This accelerated convergence indicates that the dynamic reweighting mechanism in DFT leads to more informative gradient updates, guiding the model toward high-quality solutions early in training. It also suggests that DFT helps avoid the optimization plateaus or noise-prone regions often encountered in standard SFT, thereby enabling more efficient acquisition of complex mathematical reasoning patterns.
@@
DFT  outperforms the concurrent Importance-Weighted SFT (iw-SFT) in most settings across model families and benchmarks. As shown in Table 2, DFT achieves higher average accuracy than iw-SFT on most model families: LLaMA-3.2-3B (+2.39), LLaMA-3.1-8B (+4.15), DeepSeekMath-7B (+3.34), and Qwen2.5-Math-1.5B (+1.30). Although iw-SFT slightly outperforms our method on Qwen2.5-Math-7B (+2.45), this improvement is not consistent across datasets. In particular, on the LLaMA model family, iw-SFT exhibits signs of limited robustness. For LLaMA-3.2-3B, iw-SFT underperforms standard SFT on Math500 (5.13 vs. 8.65) and AMC23 (2.03 vs. 3.13). Similarly, for LLaMA-3.1-8B, iw-SFT results in worse performance than SFT on Minerva Math (4.31 vs. 5.78) and AMC23 (7.34 vs. 8.28). These cases demonstrate that iw-SFT might struggle to generalize beyond specific training signals, and may even degrade performance under distribution shifts or on harder benchmarks. In contrast, DFT consistently improves upon both the base model and SFT across nearly all datasets, including those where iw-SFT fails. These results underline better generalization ability of DFT in diverse mathematical reasoning scenarios. Moreover, iw-SFT incurs additional computational overhead by requiring a separate reference model to compute importance weights, whereas DFTdynamically derives its own weighting directly from the model’s token probabilities, resulting in a more efficient training procedure.
4.2 Exploratory Experiment - Offline RL Setting
4.2.1 Setup and Implementation Details
Data Preparation.
We conduct an exploratory investigation of applying DFT  an offline RL setting, where the sparsity of reward issue could be alleviated comparing to SFT setting. Specifically, we adopt the commonly used rejection sampling fine-tuning (RFT) framework Dong et al. (2023); Ahn et al. (2024). Following the setup in Section 4.1, we sample responses for 10,000 math questions using a temperature of 1.0 and generate four responses per question from the base model itself. Correct responses are identified using math verify and retained as training data, resulting in approximately 140,000 examples. For DPO training, we construct 100,000 positive–negative preference pairs from the generated responses.
Training Details.
All experiments are conducted using the Qwen2.5-math-1.5B model. We compare DFT with representative offline RL methods, including DPO (Rafailov et al., 2023) and RFT (Dong et al., 2023; Ahn et al., 2024), as well as online RL methods PPO (Schulman et al., 2017) and GRPO (Shao et al., 2024). For RFT and DFT, the training setup follows the configuration in Section 4.1. For DPO, we use the ms-swift framework (Zhao et al., 2024) with a learning rate of $1\times 10^{-6}$, batch size of 128, and a warmup ratio of 0.05. For PPO and GRPO, training is performed using the verl framework (Sheng et al., 2025) with a learning rate of $1\times 10^{-6}$, batch size of 256, and a warmup ratio of 0.1. We set the number of response per input to $n=4$ for GRPO.
4.2.2 Results
Table 3: Evaluation results on five mathematical reasoning benchmarks in an offline reinforcement learning setting using reward signals from rejection sampling. DFT achieves the best overall performance, surpassing both offline (RFT, DPO) and online (PPO, GRPO) baselines, demonstrating its efficiency and strength as a simple yet effective fine-tuning strategy.
@@ 4.2.2 Results
Qwen2.5-Math-1.5B w/DPO Offline 46.89 11.53 22.86 4.58 30.16 23.20
Qwen2.5-Math-1.5B w/RFT Offline 48.23 14.19 22.29 4.37 30.78 23.97
Qwen2.5-Math-1.5B w/PPO Online 56.10 15.41 26.33 7.50 37.97 28.66
Qwen2.5-Math-1.5B w/GRPO Online 62.86 18.93 28.62 8.34 41.25 32.00
@@
DFT demonstrates best performance in the offline reinforcement learning setting, outperforming both offline and online RL baselines. As shown in Table 3, DFT achieves an average score of 35.43, exceeding the best offline method RFT by +11.46 points (from 23.97 to 35.43), and even outperforming the strongest online RL algorithm GRPO by +3.43 points (from 32.00 to 35.43). Our model performs well across all five benchmarks. On Math500, DFT scores 64.71, slightly ahead of GRPO (62.86) and better than PPO (56.10) and RFT (48.23). The gains are notable on more challenging benchmarks: on AMC23, DFT achieves 48.44, a +7.19 point margin over GRPO and a +17.66 point gain over RFT. Similarly, on Minerva Math, DFT reaches 25.16, outperforming GRPO (18.93) by +6.23 points, PPO (15.41) by +9.75, and all offline baselines by a wider gap.
@@
We also compare against the concurrent iw-SFT (Qin & Springenberg, 2025) method under the offline setting. While iw-SFT performs competitively on certain datasets, achieving 60.80 on Math500 and 44.21 on AMC23, its overall average performance (31.86) still falls short of our method by a +3.57 point margin. Moreover, iw-SFT yields only marginal improvements compared to its own performance under the standard SFT setting, achieving an average score of 31.86 in the offline RL setting versus 30.28 with SFT. This modest gain of +1.58 points stands in stark contrast to the improvement achieved by DFT (+4.76, from 30.67 to 35.43). These results suggest that iw-SFT struggles to effectively leverage reward supervision under offline constraints, whereas DFT is able to consistently convert such signals into more robust generalization and higher task performance.
4.3 Ablation and Investigation
4.3.1 Token Probability Distribution
Figure 2: Token probability distributions on the training set before training and after fine-tuning with DFT, SFT, and various RL methods including DPO, PPO, and GRPO. A logarithmic scale is used on the y-axis to improve visualization clarity.
To understand how the model trained by DFT is different from standard SFT and other RL methods, we look into the token probability distribution of the model’s output over the training set in Figure 2. The results reveal how the methods alter the probability landscape. SFT tends to uniformly increase token probabilities, shifting the entire distribution towards higher confidence, but mainly targeting the lower and lowest probability tokens. The highest probability token portion barely increases. In stark contrast, DFT exhibits a polarizing effect: it significantly boosts the probabilities of a subset of tokens while actively suppressing the probabilities of others. This leads to a bimodal distribution, with more tokens occupying both the highest and lowest probability bins. Other RL methods such as DPO, GPPO and PPO show the same trend as DFT, although the scale is much milder than it. We look into the words that belong to the lowest probability bin, and find that they are generally the conjuncative words or punctuations such as ‘the’, ‘let’, ‘,’, ‘.’ etc.
These results suggest that for robust learning, models should not attempt to fit all tokens with uniform confidence. For large language models, it may be beneficial to deprioritize fitting tokens that serve grammatical functions rather than carrying primary semantic content. This concept is analogous to human pedagogy, where students are taught to focus on substantive concepts rather than perfecting the usage of common connective words.

### v2

@@ 4.1.2 Main Results
DFT demonstrates generalization and robustness, especially on challenging benchmarks where standard SFT yields minimal or even negative impact. For instance, on Olympiad Bench, SFT degrades performance for Qwen2.5-Math-1.5B, dropping accuracy from 15.88 to 12.63, while DFT boosts it to 27.08, +11.20 point improvement over base model. On AIME24, SFT reduces accuracy for Qwen2.5-Math-7B by 4.20 points (from 6.68 to 2.48), whereas DFT improves performance to 8.56, achieving a +1.88 point gain over the base model despite the difficulty of the benchmark. A similar trend is observed on AMC23. SFT reduces the performance of Qwen2.5-Math-1.5B from 19.38 to 18.75, while DFT raises it to 38.13, a +18.75 point gain over base. For Qwen2.5-Math-7B, SFT yields only a marginal improvement (+1.86), whereas DFT achieves a +17.04 point gain. These results underscore that DFT not only scales more effectively across models of varying capacities, but also exhibits better resilience on difficult reasoning tasks where traditional SFT struggles.
DFT exhibits better learning efficiency and faster convergence characteristics. Figure 1 reveals clear differences in learning dynamics between DFT and standard SFT on Qwen2.5-Math-1.5B across all math reasoning benchmarks. Compared to SFT, our method demonstrates three distinct advantages: (1) Faster convergence, achieving peak performance within the first 120 training steps on most benchmarks; (2) Better early-stage performance, with DFT already outperforming best final accuracy of SFT within the first 10–20 steps; and (3) Higher sample efficiency, consistently requiring fewer updates to reach relatively optimal results. This accelerated convergence shows that the dynamic reweighting mechanism in DFT leads to more informative gradient updates, guiding the model toward high-quality solutions early in training. It also suggests that DFT helps avoid the optimization plateaus or noise-prone regions often encountered in standard SFT, thereby enabling more efficient acquisition of complex mathematical reasoning patterns.
We also report the results of parameter-efficient fine-tuning (PEFT) training setting (Hu et al., 2022) and training on the OpenR1-Math dataset (Hugging Face, 2025) with better quality in Appendix A.7 and Appendix A.6, respectively. Comparison and Discussion with the concurrent method iw-SFT (Qin & Springenberg, 2025) is provided in Appendix A.5.
@@
4.2 Exploratory Experiment - Offline RL Setting
Equation 7 shows that SFT suffers from reward sparsity, since in a constructed dataset each query $x$ has only a single reference answer $y^{\star}$. From the perspective of RL, RFT/RAFT (Dong et al., 2023; Ahn et al., 2024) can be viewed as alleviating the sparse reward issue by effectively increasing reward density, thereby enhancing model performance. Motivated by this observation, we conduct an exploratory study applying DFT in an offline RL setting, where the reward sparsity problem is inherently less severe compared to standard SFT, to further validate the effectiveness.
4.2.1 Setup and Implementation Details
We sample responses for 100,000 math questions using a temperature of 1.0 and generate four responses per question from the base model itself. Correct responses are identified using math verify and retained as training data, resulting in approximately 140,000 examples. For DPO training, we construct 100,000 positive–negative preference pairs from the generated responses.
We compare DFT with representative offline RL methods, including DPO (Rafailov et al., 2023) and RFT (Dong et al., 2023; Ahn et al., 2024), as well as online RL methods PPO (Schulman et al., 2017) and GRPO (Shao et al., 2024). For RFT and DFT, the training setup follows the configuration in Section 4.1. For DPO, we use the ms-swift (Zhao et al., 2024) with a learning rate of $1\times 10^{-6}$, batch size of 128, and a warmup ratio of 0.05. For PPO and GRPO, training is performed using the verl (Sheng et al., 2025) with a learning rate of $1\times 10^{-6}$, batch size of 256, and a warmup ratio of 0.1. We set the number of response $n=4$ for GRPO.
4.2.2 Results
Table 2: Evaluation results on mathematical reasoning benchmarks in an offline reinforcement learning setting using reward signals from rejection sampling. The best performance is in bold.
Setting Math500 Minerva Math Olympiad Bench AIME24 AMC23 Avg.
Qwen2.5-Math-1.5B 31.66 8.51 15.88 4.16 19.38 15.92
Qwen2.5-Math-1.5B w/DPO Offline 46.89 11.53 22.86 4.58 30.16 23.20
Qwen2.5-Math-1.5B w/RFT Offline 48.23 14.19 22.29 4.37 30.78 23.97
Qwen2.5-Math-1.5B w/PPO Online 56.10 15.41 26.33 7.50 37.97 28.66
Qwen2.5-Math-1.5B w/GRPO Online 62.86 18.93 28.62 8.34 41.25 32.00
Qwen2.5-Math-1.5B w/DFT Offline 64.71 25.16 30.93 7.93 48.44 35.43
DFT demonstrates competitive performance in the offline RL setting, outperforming both offline and online RL baselines. Table 2 shows DFT achieves an average score of 35.43, exceeding the best offline method RFT by +11.46 points, and even outperforming the strongest online RL algorithm GRPO by +3.43 points. Specially, on Math500, DFT scores 64.71, slightly ahead of GRPO (62.86) and better than PPO (56.10) and RFT (48.23). The gains are also notable on more challenging benchmarks: on AMC23, DFT achieves 48.44, a +7.19 point margin over GRPO and a +17.66 point gain over RFT. Similarly, on Minerva Math, DFT reaches 25.16, outperforming GRPO by +6.23 points, PPO by +9.75, and all offline baseline methods.
4.3 Exploratory Experiment - Code Generation Task
To further validate the effectiveness of DFT, we investigate its performance on another reasoning-intensive task: code generation.
4.3.1 Setup and Implementation Details
We adopt UltraFeedback (Cui et al., 2024) as the training dataset. From this corpus, we sample 10,000 prompts and, for each prompt, select the response with the highest average score to perform supervised fine-tuning (SFT) (Du et al., 2025). Model performance is assessed on three widely used code generation benchmarks: HumanEval (Chen et al., 2021), HumanEval+ (Liu et al., 2023), and MultiPL-E (Cassano et al., 2023). Training is conducted for one epoch with a learning rate of $5\times 10^{-5}$, a warm-up ratio of 0.05, and a batch size of 16.
4.3.2 Results
Table 3: Performance of various models on code generation benchmarks. The best performance for each benchmark is highlighted in bold.
HumanEval MultiPL-E
HumanEval HumanEval+ Python C++ Java PHP TS C# Bash JS Avg.
Qwen2.5-3B 43.3 36.0 43.29 40.99 37.34 37.89 47.17 43.04 24.68 45.96 40.05
Qwen2.5-3B w/SFT 41.5 34.8 42.07 42.24 37.97 37.27 43.40 41.77 20.25 47.83 39.10
Qwen2.5-3B w/DFT 45.7 39.0 45.73 44.72 41.77 45.34 42.14 43.04 27.85 44.10 41.84
Qwen2.5-Coder-3B 52.4 42.7 51.83 53.42 46.20 47.20 54.09 55.06 25.32 54.04 48.39
Qwen2.5-Coder-3B w/SFT 51.8 43.9 51.22 51.55 48.10 54.66 59.12 51.27 34.18 54.04 50.52
Qwen2.5-Coder-3B w/DFT 56.7 50.0 57.32 54.66 51.27 58.39 58.49 60.76 31.01 53.42 53.16
Qwen2.5-Coder-7B 62.2 53.0 63.41 63.98 53.16 59.01 62.89 59.49 39.24 60.87 57.76
Qwen2.5-Coder-7B w/SFT 54.9 48.8 54.88 64.60 51.27 62.11 68.55 60.76 33.54 65.22 57.62
Qwen2.5-Coder-7B w/DFT 67.7 59.8 67.68 67.70 54.43 60.87 70.44 65.19 48.73 63.35 62.30
Table 3 shows DFT achieves improvements in most cases compared to both base models and SFT. For Qwen2.5-3B, DFT raises HumanEval from 43.3 to 45.7 and HumanEval+ from 36.0 to 39.0, with the MultiPL-E average also increasing from 40.05 (base) and 39.10 (SFT) to 41.84. Similar trends are observed for Qwen2.5-Coder-3B, where DFT improves HumanEval to 56.7 and HumanEval+ to 50.0, outperforming both base and SFT. For Qwen2.5-Coder-7B, DFT reaches 67.7 on HumanEval, 59.8 on HumanEval+, and 62.3 average on MultiPL-E, surpassing SFT by +12.8, +11.0, and +4.7 points respectively. The overall trend demonstrates that DFT generally provides stronger performance across different models and languages.
4.4 Exploratory Experiment - Multi-Modal Reasoning
To further investigate the generalization of DFT, we extend our evaluation to multi-modal reasoning tasks that require both visual understanding and mathematical problem solving.
Table 4: Performance comparison across different multi-modal reasoning benchmarks. The best performance on each benchmark is highlighted in bold.
MathVerse MathVision WeMath
Vision Only Vision Intensive Vision Dominant Overall
Qwen2.5-VL-3B 28.81 30.96 31.60 33.83 21.25 4.10
Qwen2.5-VL-3B w/SFT 30.96 33.63 32.74 35.66 21.02 23.33
Qwen2.5-VL-3B w/DFT 32.49 35.91 33.50 37.54 22.30 23.71
4.4.1 Setup and Implementation Details
We use the WeThink dataset (Yang et al., 2025) for training. The model is fine-tuned using LLaMA-Factory (Zheng et al., 2024) and evaluated with VLMEvalKit (Duan et al., 2024). We train the model for 1 epoch with a learning rate of 5e-5. To comprehensively assess reasoning capabilities, we adopt a suite of multi-modal reasoning benchmarks including MathVerse (Zhang et al., 2024a), MathVision (Wang et al., 2024), and WeMath (Qiao et al., 2024) for evaluation.
4.4.2 Results
DFT achieves consistent improvements over base models and SFT across all multi-modal reasoning benchmarks. Table 4 shows, on MathVerse, DFT boosts Qwen2.5-VL-3B from 33.83 to 37.54 average accuracy, outperforming the SFT gain of only +1.83 by +3.71 points. Consistent improvements are observed across all major vision-related subcategories. On MathVision, DFT improves performance from 21.25 (base) to 22.30, exceeding SFT which fails to provide gains (21.02). On WeMath, SFT already yields a +19.23 point gain, but DFT pushes performance slightly further to 23.71, maintaining superiority over both base and SFT. These results indicate that DFT not only strengthens text-only reasoning but also extends effectively to multi-modal domains.
4.5 Analysis
Figure 2: Token probability distributions on the training set before training and after fine-tuning with DFT, SFT, and various RL methods. A logarithmic scale is used on the y-axis for clarity.
To understand how the model trained by DFT is different from SFT and other RL methods, we look into the token probability distribution of the model’s output over the training set in Figure 2. SFT tends to uniformly increase token probabilities, shifting the entire distribution towards higher confidence, but mainly targeting the lower and lowest probability tokens. The highest probability token portion barely increases. In stark contrast, DFT exhibits a polarizing effect: it significantly boosts the probabilities of a subset of tokens while actively suppressing the probabilities of others. This leads to a bimodal distribution, with more tokens occupying both the highest and lowest probability bins. Other RL methods such as DPO, GPPO and PPO show the same trend as DFT, although the scale is much milder than it. We look into the words that belong to the lowest probability bin, and find that they are generally the conjunctive words or punctuations such as ‘the’, ‘let’, ‘,’, ‘.’ etc. These results suggest that for robust learning, models should not attempt to fit all tokens with uniform confidence. It may be beneficial to deprioritize fitting tokens that serve grammatical functions rather than carrying primary semantic content. This concept is analogous to human pedagogy, where students are taught to focus on substantive concepts rather than perfecting the usage of common connective words. Further analysis can be found in Appendix A.4.
5 Conclusion
In this work, we address the well-documented generalization gap between Supervised Fine-Tuning and Reinforcement Learning. We provide a novel theoretical analysis, demonstrating that the standard SFT gradient is equivalent to a policy gradient update with an ill-posed implicit reward that is inversely proportional to the model’s confidence. This insight explains SFT’s tendency to overfit and its unstable optimization dynamics. Building on this analysis, we introduce DFT, a simple yet powerful method that rectifies this issue by dynamically reweighting the SFT loss with the token probability. This one-line modification stabilizes the learning process and promotes better generalization. Our extensive experiments show that DFT consistently outperforms standard SFT across various models and challenging mathematical reasoning benchmarks. Furthermore, when adapted to an offline RL setting, DFT  surpasses established online and offline RL algorithms, highlighting its effectiveness and efficiency. Our work provides both a deeper understanding of SFT and a practical, high-impact solution that significantly closes the performance gap with more complex RL methods.
Appendix A Appendix
A.1 Usage of LLM
We employ LLM primarily as writing assistants to refine and polish the manuscript. Their usage was limited to improving clarity, coherence, and presentation, while all conceptual and experimental contributions remain original.
A.2 Limitations
While our experiments demonstrate the effectiveness of DFT on mathematical reasoning benchmarks and code generation tasks, the evaluation scope remains limited. We have not yet assessed its performance on broader task categories or with larger-scale LLM, which we leave for future exploration. Moreover, DFT may not offer universal benefits across all scenarios. In domains that primarily involve the acquisition of factual knowledge, conventional SFT may remain the most efficient approach. Our aim is not to assert that DFT universally outperforms SFT, but rather to offer a new perspective on objective design by analyzing the distinction between RL and SFT.
A.3 Detailed Derivation of Equation (5)
We start from the SFT gradient in Equation (2):
$\nabla_{\theta}\mathcal{L}_{\text{SFT}}(\theta)=\mathbb{E}_{(x,y^{\star})\sim\mathcal{D}}\big[-\nabla_{\theta}\log\pi_{\theta}(y^{\star}\mid x)\big].$ (1)
For each query $x$, the expectation over expert demonstrations $(x,y^{\star})$ can be written explicitly as a summation over all possible outputs $y$:
$\mathbb{E}_{(x,y^{\star})\sim\mathcal{D}}\big[-\nabla_{\theta}\log\pi_{\theta}(y^{\star}\mid x)\big]=\mathbb{E}_{x\sim\mathcal{D}_{x}}\sum_{y}\mathbf{1}[y=y^{\star}]]\,\big[-\nabla_{\theta}\log\pi_{\theta}(y\mid x)\big].$ (2)
We insert the model distribution $\pi_{\theta}(y\mid x)$, which allows us to express the summation in terms of importance weights:
$\mathbb{E}_{x\sim\mathcal{D}_{x}}\sum_{y}\pi_{\theta}(y\mid x)\cdot\frac{\mathbf{1}[y=y^{\star}]]}{\pi_{\theta}(y\mid x)}\,\big[-\nabla_{\theta}\log\pi_{\theta}(y\mid x)\big].$ (3)
Here, the term $\tfrac{\mathbf{1}[y=y^{\star}]]}{\pi_{\theta}(y\mid x)}$ serves as an importance weight comparing the expert (Dirac delta) distribution with the model’s distribution.
The summation over $y$ can now be rewritten as an expectation under the policy distribution $y\sim\pi_{\theta}(\cdot\mid x)$:
$\mathbb{E}_{x\sim\mathcal{D}_{x}}\,\mathbb{E}_{y\sim\pi_{\theta}(\cdot\mid x)}\left[\frac{\mathbf{1}[y=y^{\star}]]}{\pi_{\theta}(y\mid x)}\big(-\nabla_{\theta}\log\pi_{\theta}(y\mid x)\big)\right].$ (4)
Thus, we obtain Equation (5):
$\mathbb{E}_{(x,y^{\star})\sim\mathcal{D}}\big[-\nabla_{\theta}\log\pi_{\theta}(y^{\star}\mid x)\big]=\mathbb{E}_{x\sim\mathcal{D}_{x}}\,\mathbb{E}_{y\sim\pi_{\theta}(\cdot\mid x)}\left[\frac{\mathbf{1}[y=y^{\star}]]}{\pi_{\theta}(y\mid x)}\big(-\nabla_{\theta}\log\pi_{\theta}(y\mid x)\big)\right].$ (5)
This derivation shows that the SFT gradient can be expressed as an on-policy policy gradient with importance sampling, where the expert demonstration distribution is reweighted relative to the model distribution.
A.4 Discussions and Insights
Gradient Analysis of DFT.
We now analyze the gradient induced by the DFT surrogate loss. Recall the sequence-level definition:
$\mathcal{L}_{\text{DFT}}(\theta)=-\,\operatorname{sg}\!\big(\pi_{\theta}(y^{\star}\mid x)\big)\;\log\pi_{\theta}(y^{\star}\mid x),$ (10)
where $\operatorname{sg}(\cdot)$ denotes the stop-gradient operator. Since the stop-gradient blocks backpropagation, the detached probability $\operatorname{sg}(\pi_{\theta}(y^{\star}\mid x))$ is treated as a constant during differentiation. Consequently, the gradient becomes
$\nabla_{\theta}\mathcal{L}_{\text{DFT}}$ $=-\,\operatorname{sg}\!\big(\pi_{\theta}(y^{\star}\mid x)\big)\;\frac{1}{\pi_{\theta}(y^{\star}\mid x)}\,\nabla_{\theta}\pi_{\theta}(y^{\star}\mid x)$ (11)
$=-\,\Bigl(\tfrac{\operatorname{sg}(\pi_{\theta}(y^{\star}\mid x))}{\pi_{\theta}(y^{\star}\mid x)}\Bigr)\,\nabla_{\theta}\pi_{\theta}(y^{\star}\mid x).$ (12)
Since $\operatorname{sg}(\pi_{\theta}(y^{\star}\mid x))$ equals $\pi_{\theta}(y^{\star}\mid x)$ in the forward pass, the prefactor is numerically equal to $1$. Therefore,
$\nabla_{\theta}\mathcal{L}_{\text{DFT}}=-\,\nabla_{\theta}\pi_{\theta}(y^{\star}\mid x).$ (13)
This shows that DFT is mathematically equivalent to directly maximizing the model probability of the target token, rather than its log-probability as in cross-entropy.
For cross-entropy, the loss is
$\mathcal{L}_{\text{CE}}(\theta)=-\log\pi_{\theta}(y^{\star}\mid x),$
yielding gradient
$\nabla_{\theta}\mathcal{L}_{\text{CE}}=-\tfrac{1}{\pi_{\theta}(y^{\star}\mid x)}\nabla_{\theta}\pi_{\theta}(y^{\star}\mid x).$
Thus both CE and DFT share the same gradient direction but differ in scaling: CE amplifies updates for low-probability targets (factor $1/\pi$), while DFT applies a uniform factor $1$. As a result, DFT avoids the instability caused by excessively large gradients on unlikely expert tokens, providing more conservative and stable updates.
From the reinforcement learning perspective, the reward of DFT becomes uniformly $1$ across all expert trajectories, equivalent to a verification-style objective that treats all correct references equally. From the optimization perspective, DFT trades off aggressive fitting of rare tokens for better stability and calibration. Practically, this explains why DFT often yields smoother training and stronger generalization, while maintaining alignment with the pre-training distribution.
Learning from Noisy Data.
DFT offers a simple yet effective approach, prompting us to reflect on why it might actually work. One intuitive explanation lies in its ability to learn from noisy data (Freund, 2009). Sasaki & Yamashina (2020) propose an imitation learning algorithm for learning from noisy demonstrations, based on the core idea of avoiding the fitting of data that is difficult to model, as such data is likely to originate from suboptimal behaviors, i.e., noise. Their method introduces a weighted behavioral cloning objective, where the weights are derived from a previously trained policy’s confidence in each action. Similarly, the weighting mechanism in DFT  shares the same intuition, but instead of relying on a fixed old policy model to compute confidence scores, it uses a single policy model to perform confidence-based weighting on-the-fly during training.
A.5 Comparision with Concurrent work iw-SFT
We include a concurrent method, Importance-Weighted SFT (iw-SFT) (Qin & Springenberg, 2025), for comparison. All training settings follow those reported in the original paper, except that we set the number of training epochs to 1.
As shown in Table 5, DFT achieves higher average accuracy than iw-SFT on most model families: LLaMA-3.2-3B (+2.39), LLaMA-3.1-8B (+4.15), DeepSeekMath-7B (+3.34), and Qwen2.5-Math-1.5B (+1.30). Although iw-SFT outperforms our method on Qwen2.5-Math-7B (+2.45), this improvement is not consistent across datasets. In particular, for LLaMA-3.2-3B, iw-SFT underperforms standard SFT on Math500 (5.13 vs. 8.65) and AMC23 (2.03 vs. 3.13). Similarly, for LLaMA-3.1-8B, iw-SFT results in worse performance than SFT on Minerva Math (4.31 vs. 5.78) and AMC23 (7.34 vs. 8.28). In contrast, DFT consistently improves upon both the base model and SFT across nearly all datasets, including those where iw-SFT fails. These results underline better generalization ability of DFT in diverse mathematical reasoning scenarios. Moreover, iw-SFT incurs additional computational overhead by requiring a separate reference model to compute importance weights, whereas DFT dynamically derives its own weighting directly from the token probabilities of model, resulting in a more efficient training procedure.
@@ A.5 Comparision with Concurrent work iw-SFT
Table 6: Evaluation results on five mathematical reasoning benchmarks in the offline reinforcement learning setting using rejection-sampling–based reward signals, compared against iw-SFT.
@@
We also compare against iw-SFT under the offline setting, as shown in Table 6. While iw-SFT performs competitively on certain datasets, achieving 60.80 on Math500 and 44.21 on AMC23, its overall average performance (31.86) remains below that of our method by +3.57 points. Moreover, iw-SFT shows only modest improvements compared to its standard SFT counterpart, with an average score of 31.86 in the offline RL setting versus 30.28 with SFT (+1.58). In contrast, DFT achieves a larger gain of +4.76 (from 30.67 to 35.43). These results indicate that iw-SFT provides limited benefits from reward supervision under offline constraints, whereas DFT is able to more effectively incorporate such signals, leading to better generalization and higher task performance.
A.6 Exploratory Experiment - OpenR1-Math Training Dataset
Inspired by DeepSeek-R1 DeepSeek-AI et al. (2025), several studies have attempted to train open-source models to reproduce its reasoning capabilities (Hugging Face, 2025). To this end, a high-quality dataset, OpenR1-Math-220k (Hugging Face, 2025), was constructed, where the prompts are drawn from NuminaMath 1.5 and the off-policy reasoning traces are generated by DeepSeek-R1. LUFFY (Yan et al., 2025) further filtered out sequences longer than 8192 tokens as well as those verified incorrect by Math-Verify, resulting in about 45k prompts paired with off-policy reasoning traces. We adopt this dataset as the training corpus for SFT. All training details remain the same as previous experiments, except that the number of epochs is set to 3.
As shown in Table 7, training on OpenR1-Math-220k consistently improves performance, and the use of higher-quality annotations yields additional gains. SFT on this dataset increases the average accuracy of Qwen2.5-Math-1.5B by +13.24 points compared to the base model, while DFT provides a further +9.03 gain, resulting in a total improvement of +22.27. These results suggest that DFT remains effective even when applied on top of high-quality training data, highlighting its potential as a general fine-tuning paradigm.
Table 7: Performance training on the OpenR1-Math-220k dataset. The best score for each benchmark is in bold.
Math500 Minerva Math Olympiad Bench AIME24 AMC23 Avg.
Qwen2.5-Math-1.5B 31.66 8.51 15.88 4.16 19.38 15.92
Qwen2.5-Math-1.5B w/SFT 61.60 20.29 24.27 4.16 35.47 29.16
Qwen2.5-Math-1.5B w/DFT 71.76 27.00 33.48 9.79 48.91 38.19
Table 8: Performance on five mathematical reasoning benchmarks using LoRA for training. The best score of each model across benchmarks is highlighted in bold.
Math500 Minerva Math Olympiad Bench AIME24 AMC23 Avg.
LLaMA-3.2-3B 1.63 1.36 1.01 0.41 1.56 1.19
LLaMA-3.2-3B w/SFT 4.88 1.56 1.68 0.00 2.66 2.56
LLaMA-3.2-3B w/DFT 11.13 5.18 3.87 0.00 2.97 4.63
Qwen2.5-Math-1.5B 31.66 8.51 15.88 4.16 19.38 15.92
Qwen2.5-Math-1.5B w/SFT 41.47 10.85 11.56 1.45 17.03 16.87
Qwen2.5-Math-1.5B w/DFT 64.85 22.58 28.45 5.84 40.78 32.90
A.7 Exploratory Experiment - PEFT Training Setting
To investigate whether DFT remains effective under parameter-efficient fine-tuning (PEFT) settings with limited compute, we apply DFT using LoRA adapters across two model families: LLaMA-3.2-3B and Qwen2.5-Math-1.5B. All training configurations remain identical to previous full-parameter experiments, except that LoRA is enabled with rank=8 and alpha=16.
As shown in Table 8, DFT provides consistent improvements over both base and SFT baselines under LoRA-based PEFT. For Qwen2.5-Math-1.5B, DFT increases the average accuracy from 15.92 (base) and 16.87 (SFT) to 32.90. For LLaMA-3.2-3B, DFT achieves a gain of +3.44 over SFT (from 1.19 to 4.63). These results indicate that DFT can serve as an effective fine-tuning strategy in low-resource or compute-constrained settings, where full model updates are not practical.
A.8 Training Hyper-Parameters Ablation